# Enhancing Temporal Consistency in Video Editing by Reconstructing Videos with 3D Gaussian Splatting

### v3

@@ D.1 Video-3DGS (1st Stage)
2D Learnable Parameter for Merging Foreground and Background Views We visualized the 2D learnable parameter (i.e., $\alpha$ in Equation 5) in different video scenarios in the following link: https://video-3dgs-2d-learnable-mask.github.io/. As shown in the videos, the learned $\alpha$ focuses on the foreground regions, while the learned (1 - $\alpha$) focuses on the background regions.
@@ D.3 Relationship between Video-3DGS (1st Stage) and Video-3DGS (2nd Stage)
We examine the correlation between video reconstruction and editing quality in Table 8. We introduce two versions of Video-3DGS: model A, where all clips utilize the same deformation field (thus reducing the representation capacity), and model B, our default setting. As depicted in the table, model B achieves superior reconstruction quality compared to model A, resulting in better performance across all three off-the-shelf video editors. We also conducted qualitative comparison with 3DGS (Kerbl et al., 2023) to further prove the positive relationship between video reconstruction and video editing in https://video-3dgs-recon-edit.github.io/.
@@ E.1 Video Reconstruction
For video reconstruction, we visualize qualitative results in Figure 7, Figure 8, and Figure 9. As shown in the figures, the proposed Video-3DGS consistently demonstrates higher reconstruction quality. Video comparisons with 3DGS (Kerbl et al., 2023) and Deformable-3DGS (Yang et al., 2023) are provided in the following link: https://video-3dgs-additional-recon.github.io/
@@ E.1.1 Spatial Decomposition
We also provide visual analysis on the spatial decomposition in the following video link: https://video-3dgs-spatial.github.io/. Our method successfully extracts the corresponding 3D points of moving objects using MC-COLMAP. We also found that the spatial decomposition can work well in a scene with multiple foreground objects.
@@ E.2.1 Single-phase Refiner
To show the visual effectiveness of Video-3DGS as single-phase refiner, we visualize qualitative results in Figure 10Figure 11, and Figure 12, which adopt Text2Video-Zero (Khachatryan et al., 2023), TokenFlow (Geyer et al., 2023), and RAVE (Kara et al., 2023) as the underlying video editor, respectively. As shown in the figures, Video-3DGS effectively enhances the temporal consistency in the edited results across all three video editors. We further provide the video analysis of Video-3DGS as a plug-and-play refiner to show better visual comparison in this link: https://video-3dgs-additional-edit.github.io/.
@@ E.2.2 Recursive and Ensembled Refiner
We further conduct the visual analysis on Video-3DGS as Recursive and Ensembled (RE) refiner. First, similar to Figure 3, we provide more video results showing the sensitivity of video editors to hyperparameter changes in this link: https://video-3dgs-re-hyperparameters.github.io/. It confirms that different hyperparameter values in the deployed video editors results in different outputs. Then, we show the editing visual improvement brought by Video-3DGS (RE) in https://video-3dgs-re.github.io/.
@@ Appendix F Asset Licenses
Appendix G Discussion & Limitations
While the proposed Video-3DGS exhibits remarkable performance in both video reconstruction and editing tasks, certain challenges persist. In video reconstruction, faithful reconstruction becomes challenging when foreground objects exhibit extremely large motion. In video editing, unsatisfactory results may arise if the underlying video editor fails or if significant changes to object shapes are required, which contradicts the inherent property of 3D Gaussians prioritizing the preservation of original structure. Nevertheless, our work represents a pioneering effort in extending 3DGS to enhance the temporal consistency of initially edited videos from any zero-shot video editor. We envision that this will inspire further research in a zero-shot video editing community. Unlike other conventional neural representation papers, our work does not explore the task of novel view synthesis. Instead, our method focuses on video reconstruction and editing by encoding and fitting videos with 3DGS. We anticipate that our Video-3DGS will also serve as a fundamental framework for 4D novel view synthesis in future research.
Appendix H Broader Impact
@@
Figure 10: Qualitative results of single-phase refiner for video editing by adopting Text2Video-Zero.
Figure 11: Qualitative results of single-phase refiner video editing by adopting TokenFlow.
Figure 12: Qualitative results of single-phase refiner for video editing by adopting RAVE.

### v4

@@ D.1 Video-3DGS (1st Stage)
Ablations on Segmentation Modules we additionally include the ablation study on using different video segmentation modules. There are several open-vocabulary video object segmentation models available, and we have chosen the most representative models for our experiments, including DEVA model with SAM, and Mobile-SAM. We experimented with three settings: (a) DEVA with SAM and user prompts (i.e., user needs to provide the object class names), (b) DEVA with SAM and VIPSeg foreground categories (i.e., we use the foreground categories defined by VIPSeg dataset without any user prompts), and (c) DEVA with Mobile-SAM and VIPSeg foreground categories. As shown in this video, our Video-3DGS is robust to the change of segmentation modules. Furthermore, it can also prove that our MC-COLMAP effectively reconstructs foreground moving objects regardless of their predefined categories, thereby achieving comparable reconstruction capabilities.
DAVIS (480p) PSNR SSIM WarpSSIM Qedit
CoDeF 28.7 0.904 0.926 23.7
Video-3DGS (3k) 40.8 0.986 - -
CoDeF+Video-3DGS - - 0.901 23.2
RAVE - - 0.872 23.4
RAVE+Video-3DGS - - 0.908 24.1
RAVE+Video-3DGS (RE) - - 0.913 24.8
Youtube PSNR SSIM WarpSSIM Qedit
CoDeF 25.5 0.795 0.925 21.8
Video-3DGS (3k) 40.8 0.986 - -
CoDeF+Video-3DGS - - 0.924 21.2
TokenFlow - - 0.848 22.1
TokenFlow+Video-3DGS - - 0.923 23.1
TokenFlow+Video-3DGS (RE) - - 0.923 24.2
Table 6: Comparison of the proposed Video-3DGS method with the CoDEF Ouyang et al. (2023) framework in the context of video editing tasks. The results demonstrate that Video-3DGS exhibits strong compatibility with zero-shot video editors, yielding superior or comparable performance to CoDEF across several video editing benchmarks.
@@
2D Learnable Parameter for Merging Foreground and Background Views We visualized the 2D learnable parameter (i.e., $\alpha$ in Equation 5) in different video scenarios in the following link. As shown in the videos, the learned $\alpha$ focuses on the foreground regions, while the learned (1 - $\alpha$) focuses on the background regions.
@@ D.2 Video-3DGS (2nd Stage)
Comparison With NeRF-based Video Editing  In addition to Table 2 where we show the compatibility of Video-3DGS with zero-shot video editors, we compare with a training-based video editor that utilizes NeRF Mildenhall et al. (2021) CoDEF Ouyang et al. (2023) as in Table 6. We can observe that Video-3DGS with zero-shot video editors (e.g., RAVE or TokenFlow) show comparable or better results than “CoDeF+Video-3DGS” in terms of both reconstruction and editing ability. Qualitative Results with CoDEF can be found in this link. As shown in the qualitative results, CoDEF struggles to preserve the content and structure of the original video due to its application of ControlNet on a canonical image. In contrast, our Video-3DGS effectively maintains the original content and structure.
Method Metrics WarpSSIM↑ $Q_{edit}$
UniEdit (Bai et al., 2024) 0.779 19.95
+Video-3DGS 0.830 21.15
AnyV2V (Ku et al., 2024) 0.737 19.85
+Video-3DGS 0.836 22.25
Table 7: Additional results on compatibility of Video-3DGS with video editors using video generation models. Our Video-3DGS shows consistent compatibility with video editors using video generation model by increasing the editing quality. We used 16 DAVIS videos with 32 frames and initially edited them with style-change prompts from LOVEU (Wu et al., 2023b) benchmark.
Comparison With Editors Using Video Generation Models  Recently, a new trend in video editing has emerged by leveraging video generation models. These approaches can be broadly categorized into two types: those that use image-to-video (I2V) priors and those that employ text-to-video (T2V) priors. To illustrate, we selected a representative method from each category—UniEdit (Bai et al., 2024) (T2V-based) and AnyV2V (Ku et al., 2024) (I2V-based). Due to UniEdit’s high memory requirements, we divided DAVIS videos (32 frames each) into four chunks and applied the LaVie (Wang et al., 2024b) T2V model to each chunk as a prior for video editing. For AnyV2V, we edited the first frame using Instruct-pix2pix (Brooks et al., 2023) and then propagated the edits to the remaining frames using its I2V model. After obtaining the initial outputs from each methods, we refined the results with our Video-3DGS. As shown in Table 7, Video-3DGS demonstrates strong compatibility with recent methods that use video generation priors, particularly when processing DAVIS videos with style-change prompts from the LOVEU benchmark.
@@ D.3 Relationship between Video-3DGS (1st Stage) and Video-3DGS (2nd Stage)
We examine the correlation between video reconstruction and editing quality in Table 10. We introduce two versions of Video-3DGS: model A, where all clips utilize the same deformation field (thus reducing the representation capacity), and model B, our default setting. As depicted in the table, model B achieves superior reconstruction quality compared to model A, resulting in better performance across all three off-the-shelf video editors. We also conducted qualitative comparison with 3DGS (Kerbl et al., 2023) to further prove the positive relationship between video reconstruction and video editing in this link.
@@ E.1 Video Reconstruction
For video reconstruction, we visualize qualitative results in Figure 7, Figure 8, and Figure 9. As shown in the figures, the proposed Video-3DGS consistently demonstrates higher reconstruction quality. Video comparisons with 3DGS (Kerbl et al., 2023) and Deformable-3DGS (Yang et al., 2023) are provided in the following link
@@ E.1.1 Spatial Decomposition
We also provide visual analysis on the spatial decomposition in the following video link. Our method successfully extracts the corresponding 3D points of moving objects using MC-COLMAP. We also found that the spatial decomposition can work well in a scene with multiple foreground objects.
@@ E.2.1 Single-phase Refiner
To show the visual effectiveness of Video-3DGS as single-phase refiner, we visualize qualitative results in Figure 10Figure 11, and Figure 12, which adopt Text2Video-Zero (Khachatryan et al., 2023), TokenFlow (Geyer et al., 2023), and RAVE (Kara et al., 2023) as the underlying video editor, respectively. As shown in the figures, Video-3DGS effectively enhances the temporal consistency in the edited results across all three video editors. We further provide the video analysis of Video-3DGS as a plug-and-play refiner to show better visual comparison in this link.
@@ E.2.2 Recursive and Ensembled Refiner
We further conduct the visual analysis on Video-3DGS as Recursive and Ensembled (RE) refiner. First, similar to Figure 3, we provide more video results showing the sensitivity of video editors to hyperparameter changes in this link. It confirms that different hyperparameter values in the deployed video editors results in different outputs. Then, we show the editing visual improvement brought by Video-3DGS (RE) in here.
@@ Appendix F Asset Licenses
Appendix G Broader Impact
@@
Figure 10: Qualitative results of single stage refiner for video editing by adopting Text2Video-Zero.
Figure 11: Qualitative results of single stage refiner video editing by adopting TokenFlow.
Figure 12: Qualitative results of single stage refiner for video editing by adopting RAVE.